# Periodic Chains Scheduling on Dedicated Resources - A Crucial Problem in Time-Sensitive Networks

### v1

@@ 4 Solution Approach and Algorithms
Figure 4: Outline of the proposed approach predecessor-flow.
This section describes the proposed solution approach for solving PSPs with harmonic periods. The approach is based on our local search [30], which uses the first fit scheduling (FFS) heuristic. The overall flow, visualized in Fig. 4, shows the local search and our two CP-based improvements. If the local search (Section 4.1) is unable to find a feasible solution quickly, the first improvement provides it with the warm start (Section 4.2); this is particularly useful for solving highly-utilized instances. The second improvement reduces the degeneracy of the found schedule even further using a matheuristic (Section 4.3).
@@ 4.1 Local Search
We outlined the local search method in [30]. The inner FFS heuristic processes the tasks following the ordered list of tasks $\mathcal{J}$; the ordered lists are generated and supplied to the heuristic by the encompassing local search. One by one, the tasks are scheduled on their dedicated resources as shown in Algorithm 1.
We distinguish two methods. The leftmost method schedules each task as early as possible so that none of its occurrences overlap with any occurrence of any already scheduled tasks (for more details, see Section 4 in [30]). This method is more suitable for finding a feasible solution, but it may result in a large degeneracy since it ignores the precedence constraint. The second predecessor method performs the following: If predecessor $i$ of the task $i+1$ in chain $k$ is already scheduled, we schedule the current task only after the completion time of its predecessor. Unlike the previous method, the predecessor method does not ignore the precedence of already scheduled predecessors. In both cases, the actual schedule is obtained by postponing the tasks that do not adhere to the precedence constraints.
@@
postpone tasks to satisfy precedence constraints
@@
We use local search to find the appropriate ordered list $\mathcal{J}$ to provide to the FFS heuristic. Starting with an initial ordered list, the local search first tries to modify it to respect the precedences of chains. Once this fails, the local search performs either random moves or chain reordering. The random moves are performed by swapping two tasks in the ordered list. The reordering operation takes one (arbitrary) chain whose order of tasks in the list does not follow its precedence constraints and reorders these tasks to follow them.
In the first iteration, the local search begins with a rate-monotonic ordered list, which sorts tasks in ascending order of their periods. As we discussed in [12], the rate-monotonic ordered list is the desirable starting point since we process most constrained tasks (i.e., tasks with most occurrences) at the start of the procedure when the schedule is mostly empty.
@@ 4.2 Warm-Start by Constraint Programming
The local search approach generally works very well. However, with the utilization of instances close to 100%, the approach struggles to find any feasible solution. In that case, it is beneficial to provide the local search with an alternative initial ordered list, which heuristic maps to a feasible schedule immediately.
@@ 4.2.1 Feasible Schedule for a Single Resource
We use the bin-packing CP formulation for monoprocessor (as described in [12]) relying on the height-divisible packing of Section 2.3.2. The CP solver processes one resource at a time and creates a feasible schedule for each one of them. Afterward, these feasible schedules are used to infer a suitable ordered list for the local search.
Set of tasks $\mathcal{J}_{m}$ (with periods $\tau(1),\dots,\tau(j),\dots,\tau(\rho)$) induces an associated set of rectangles $\mathcal{R}_{m}$, whose heights and widths are given as in Section 2.3.2. The set of rectangles’ heights corresponds to the periods of the tasks: $h_{j}=\tau(\rho)/\tau(j)$. Thus, $h_{1}=H$, which is equal to the height of the entire bin, and $h_{\rho}=1$.
@@
Our formulation packs the rectangles of $\mathcal{R}_{m}$ in a height-divisible manner to produce a packing in canonical form; for each rectangle it holds, that all rectangles to its left have a height greater or equal. See the example in Fig. 5, which shows canonical packing corresponding to Fig. 2(b) (if there exists a feasible packing, then there exists a feasible canonical packing - see the proof in [12]). The T-like bold lines in the figure show how the bin is recursively partitioned into sub-bins, where rectangles with smaller heights can be packed, given the height-divisible property of packing. Rectangles with height $h_{1}=H=6$ can only be packed into a single sub-bin that spans the entire figure. The rectangles with height $h_{2}=3$ have two available sub-bins. Finally, the rectangles with $h_{3}=1$ can be packed into six different sub-bins. However, since only three of these sub-bin of height 1 are filled with rectangles, the remaining three have width 0 and are not shown in Fig. 5.
To utilize this sub-bin view of the packing, we partition the rectangles to lists $\mathcal{R}_{m}^{j}$ according to their height $h_{j}$ (an example is described later in the section). Each such list is associated with a vector of widths of the rectangles $\mathbf{l}^{j}$ and a vector of integer assignment variables $\mathbf{a}^{{j}}$ with domain $\{0,\cdots,(H/h_{j})-1\}$. The assignment variable determines which sub-bin the rectangle will be packed into. Furthermore, we associate each available sub-bin $k$ of height $h_{j}$ with one variable in the load vector $\mathbf{L}^{j}=(L^{j}_{0},\dots,L^{j}_{k},\dots,L^{j}_{(H/h_{j})-1})$. The load variable keeps track of how much the sub-bin is filled. The width of the bin $w$, into which we pack the rectangles, is given by the shortest period $\tau(1)$.
@@
The pack constraint forces the model to assign each rectangle to one of the appropriate sub-bins and sets the associate load variable of the sub-bin $L^{j}_{k}$ to the sum of the widths of assigned rectangles.
For example, Fig. 5 contains five rectangles of height $h_{2}=3$ in $\mathcal{R}_{m}^{2}=(3,4,5,6,7)$ with corresponding widths given by $\mathbf{l}^{2}=(17,9,6,5,2)$. The top sub-bin for $h_{2}=3$ is occupied by rectangles 3, 4, and 7, resulting in $L^{2}_{1}=28$. The bottom sub-bin for $h_{2}=3$ is occupied by rectangles 5 and 6, resulting in $L^{2}_{0}=11$. Therefore, the vector of assignment variables is $\mathbf{a}^{{2}}=(1,1,0,0,1)$.
The second constraint enforces that for each row of height 1, the sum of loads of sub-bins that cross the row (across the different heights) does not exceed the bin’s width. In case of row 2 in Fig. 5, we have:
$L^{1}_{0}+L^{2}_{0}+L^{3}_{2}=12+11+13=36\leq 40=w$
Once the CP solver finds a feasible HD2D packing in canonical form (given the values of assignment variables $\mathbf{a}^{j}$), we transform it to the schedule of the original single resource PSP. The coordinates of the rectangles determine the start times of the associated tasks.
@@ 4.2.2 Inferring Ordered List from Schedule
An ordered list that should produce a feasible solution for leftmost heuristic is obtained simply by sorting the tasks by their start time. However, to be used with the predecessor method, this relatively arbitrary ordered list may lead to unexpected behavior since it does not respect inter-resource coupling via precedences. We attempt to resolve this by employing an order-determining CP model for the predecessor method. In this model, each task $J_{i}^{k}$ of the instance is associated with integer variable $a^{k}_{i}$. All variables $a^{k}_{i}$ are members of set $\mathbf{a}$. The CP model is:
$\mathtt{AllDiff}(\mathbf{a})$ (5)
$a^{k}_{i}<a^{l}_{j}$ $\forall\mathcal{J}_{m}\leavevmode\nobreak\ \forall J_{i}^{k},J_{j}^{l}\in\mathcal{J}_{m}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }t_{i}^{k}<t_{j}^{l}$ (6)
$a^{k}_{i+1}<a^{k}_{i}$ $\forall\mathcal{C}^{k}\leavevmode\nobreak\ \forall J_{i}^{k},J_{i+1}^{k}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }t_{i}^{k}+p_{i}^{k}>t_{i+1}^{k}$ (7)
This model processes all the tasks (and thus also the resources) simultaneously. It directly determines the ordered list of tasks with the ”all-different” constraint - the value of $a_{i}^{k}$ is the position of $J_{i}^{k}$ in the ordered list. Furthermore, when the successor task $J_{i+1}^{k}$ would be scheduled later than planned due to the start time of its predecessor, we ensure it is processed before its predecessor so the heuristic does not unexpectedly postpone it.
@@ 4.2.3 Feasibility Issues
The aforementioned CP approach consistently finds feasible solutions for 100% utilized instances. Nevertheless, when the utilization drops, it may happen that the heuristic cannot transform the provided order into a feasible solution. However, this happens at worst for 10% of instances on the most affected instance set. Furthermore, as we show in Section 5, the local search successfully finds feasible solutions for less-utilized instances even without a warm start.
@@
Figure 6: Schedules documenting the issue with a warm start on less-utilized resources. Notice the gap caused by scheduling task 123 at an incorrect time caused by FFS, leading to infeasibility.
This is demonstrated in Fig. 6. In Fig. 6(a), the schedule of one specific resource produced by the CP warm start is clearly feasible. However, when the FFS heuristic processes tasks row by row, we can see in Fig. 6(b) that it misplaces task 123; this CP model would put the task into the second row, but the heuristic schedules it in the first one, as it operates in a first fit manner, and there is excessive space available. This creates a gap due to the task’s periodicity. Finally, this inefficiency makes it impossible to schedule task 187 with a long processing time later.
@@ 4.3 Constraint Programming Matheuristic
The local search performs generally well, but it sometimes struggles to minimize the degeneracy of highly-utilized instances. We demonstrate this using Fig. 7 evaluating the local search on a set of more difficult instances. In Fig. 7(a), each curve shows the degeneracy of the best-known solution of instance as the local search progresses. We see that local search did not converge for several instances. However, in other cases, the local search rapidly decreased the degeneracy and stagnated for the rest of the experiment.
We investigated how the degeneracy is distributed within these stagnating instances. Let $\mathbf{J}_{m}^{\tau(\rho)}$ be set of tasks of machine $m$ with period $\tau(\rho)$. Similarly, let $\mathbf{C}^{\tau(\rho)}$ be sets of chains with period $\tau(\rho)$. Blue dots in Fig. 7(b) show the values of $\phi^{\tau(\rho)}=\frac{\sum_{\mathcal{C}^{k}\in\mathbf{C}^{\tau(\rho)}}D^{k}}{D_{sum}}[\%]$ for each such instance; instances are sorted in ascending order of $\phi^{\tau(\rho)}$ in the chart. The metric represents the percentage of the degeneracy left in the chains with the longest period (i.e., $T^{k}=\tau(\rho)$, the hyperperiod). To our surprise, these longest-period chains produce 40%-80% of degeneracy, even though they have just a single occurrence in the schedule.
Furthermore, by orange dots we show values of $\Omega^{\tau(\rho)}=\frac{\sum_{J_{i}^{k}\in\mathbf{J}_{m}^{\tau(\rho)}}p_{i}^{k}/T^{k}}{U_{m}}[\%]$; it shows the ratio of the utilization caused by the longest period tasks, calculated the most utilized resource $m$. We can see that this value is lower than $\phi^{\tau(\rho)}$. This suggests that while the longest-period tasks do not necessarily utilize the instance the most, their chains significantly contribute to the overall degeneracy. The culprit is probably the fact that the ordered list generated by the local search is generally almost rate-monotonic. Since the FFS heuristic processes these tasks later, there are fewer options to schedule them with low degeneracy.
(a) Degeneracy of the current best solution over time. Each line corresponds to a single instance.
(b) Ratio of degeneracy $\phi^{\tau(\rho)}$ and ratio of utilization $\Omega^{\tau(\rho)}$ covered by the longest-period chains and tasks.
@@
This observation motivated us to use a matheuristic to improve the schedule further. Similarly, as in [36, 37], we focus on a natural way to isolate part of the problem: individual periods. If we fixed all tasks with periods shorter than $\tau(\rho)$, we could try to re-schedule the longest-period tasks within the complete hyperperiod $\left[0;\tau(\rho)-1\right]$. Other tasks behave as fixed obstacles. Similarly, we can focus on optimizing chains with arbitrary period $\tau(\beta)$. With this focus on a single period, we try to optimize the schedule using a CP solver, exploiting its capabilities for solving common scheduling problems using interval variables.
@@
Let $\mathcal{J}_{m}^{O}$ be a set of tasks for the period $\tau(\beta)$ we are trying to optimize, and let $\mathcal{J}_{m}^{F}$ be a set of the tasks of other periods with their start time fixed. This is done simultaneously for all resources together but focusing on a single period only. As in Section 2.3, we shift each resource’s start time so that a task with the shortest period starts at time 0. Let $o_{m}$ be such offset. Again, let $s_{i}^{k}$ be the shifted core schedule of the task, as it is in Section 2.3. Finally, set of chains with period $T^{k}=\tau(\beta)$ is denoted as $\mathbf{C}^{\tau(\beta)}$.
@@
We formulate the following matheuristic CP model. The model tries to re-schedule tasks with period $\tau(\beta)$ to decrease the criterion while keeping the other tasks fixed:
$\min\leavevmode\nobreak\ \sum_{\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}}(\lceil S^{k}/\tau(\beta)\rceil-1)$ (8)
$\mathtt{NoOverlap}(\mathcal{J}_{m}^{O}\cup\mathcal{J}_{m}^{F})$ $\forall{m}$ (9)
$\mathtt{StartOf}(J_{i}^{k})=s_{i}^{k}$ $\forall{m}\forall J_{i}^{k}\in\mathcal{J}_{m}^{F}$ (10)
$\mathtt{StartOf}(J_{i}^{k})\geq 0$ $\forall{m}\forall J_{i}^{k}\in\mathcal{J}_{m}^{O}$ (11)
$\mathtt{EndOf}(J_{i}^{k})\leq\tau(\beta)$ $\forall{m}\forall J_{i}^{k}\in\mathcal{J}_{m}^{O}$ (12)
$ps_{i}^{k}=(\mathtt{StartOf}(J_{i}^{k})+o_{\pi_{i}^{k}})\mod\tau(\beta)$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k}\in\mathcal{C}^{k}$ (13)
$pe_{i}^{k}=(\mathtt{EndOf}(J_{i}^{k})+o_{\pi_{i}^{k}})\mod\tau(\beta)$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k}\in\mathcal{C}^{k}$ (14)
${df}_{0}^{k}=0$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}$ (15)
$ps_{i+1}^{k}-pe_{i}^{k}\geq 0\implies{df}_{i+1}^{k}=ps_{i+1}^{k}-pe_{i}^{k}$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k},J_{i+1}^{k}\in\mathcal{C}^{k}$ (16)
$ps_{i+1}^{k}-pe_{i}^{k}<0\implies{df}_{i+1}^{k}=ps_{i+1}^{k}-pe_{i}^{k}+\tau(\beta)$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k},J_{i+1}^{k}\in\mathcal{C}^{k}$ (17)
$S^{k}=\sum_{\forall J_{i}^{k}\in\mathcal{C}^{k}}p_{i}^{k}+{df}_{i}^{k}$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}$ (18)
@@
$J_{i}^{k}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathtt{intervalVar}$ $\forall{m}\forall J_{i}^{k}\in\mathcal{J}_{m}^{F}\cup\mathcal{J}_{m}^{O}$ (19)
@@
$S^{k}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathtt{integerVar}$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}$ (20)

### v2

@@ 4 Solution Approach and Algorithms
Figure 5: Outline of the proposed approach.
This section describes the proposed solution approach for solving PSPs with harmonic periods. The approach is based on local search [35], which uses the first fit scheduling (FFS) heuristic. The overall flow, visualized in Fig. 5, shows the local search and two CP-based improvements. If the local search (Section 4.1) is unable to find a feasible solution quickly, the first improvement provides it with the warm start (Section 4.2); this is particularly useful for solving highly-utilized instances. The second improvement reduces the degeneracy of the found schedule even further using a matheuristic (Section 4.3).
@@ 4.1 Local Search
The local search method is outlined in [35]. The inner FFS heuristic processes the tasks following the ordered list of tasks $\mathcal{J}$; the ordered lists are generated and supplied to the heuristic by the encompassing local search. One by one, the tasks are scheduled on their dedicated resources as shown in Algorithm 2.
We distinguish two methods. The leftmost method schedules each task as early as possible so that none of its occurrences overlap with any occurrence of any already scheduled tasks (for more details, see Section 4 in [35]). This method is more suitable for finding a feasible solution, but it may result in a large degeneracy since it ignores the precedence constraint. The second predecessor method performs the following: If predecessor $i$ of the task $i+1$ in chain $k$ is already scheduled, we schedule the current task only after the completion time of its predecessor. Unlike the previous method, the predecessor method does not ignore the precedence of already scheduled predecessors. Note that by scheduling the successor task after alternatively selected time $t_{i}^{k}+l_{i,i+1}^{k}$, we could modify the algorithm to consider both link delays ($l_{i,i+1}^{k}>p_{i}^{k}$) as well as cut-through mechanisms ($l^{k}_{i,i+1}<p_{i}^{k}$). In both cases, the actual schedule is obtained by postponing the tasks that do not adhere to the precedence constraints, according to Lemma 1.
@@
foreach $\mathcal{C}^{k}$ do
       // postpone tasks to satisfy precedences
       for $i\leftarrow 1$ to $n^{k}-1$ do
             $q\leftarrow\min q\leavevmode\nobreak\ |\leavevmode\nobreak\ t_{i+1}^{k}+q\cdot T^{k}\geq t_{i}^{k}+p_{i}^{k}$
            $t_{i+1}^{k}\leftarrow t_{i+1}^{k}+q\cdot T^{k}$
       end for
end foreach
@@
We use local search to find the appropriate ordered list $\mathcal{J}=(J_{1},\dots,J_{N})$, where $N$ is a total number of tasks, which is provided to the FFS heuristic. Given an initial ordered list, the local search first tries to modify it to respect the precedences of chains, one chain $k$ at a time, e.g.:
$(\dots,J_{2}^{k},J_{3}^{k},\dots,J_{1}^{k},\dots)\rightarrow(\dots,J_{1}^{k},J_{2}^{k},\dots,J_{3}^{k},\dots)$ (47)
Once this fails, the local search performs either random moves or chain reordering, each with 50 % probability. The random moves are performed by swapping two tasks in the ordered list.
$(\dots,J_{i}^{k},\dots,J_{j}^{l},\dots)\rightarrow(\dots,J_{j}^{l},\dots,J_{i}^{k},\dots)$ (48)
When a random move is performed, the tasks to be swapped are either two arbitrary tasks, two tasks of the same chain ($k=l)$, or two tasks that are consecutive in the given chain ($k=l\wedge j=i+1)$. The variant of the random move is selected uniformly random. The reordering operation takes one (arbitrary) chain whose order of tasks in the list does not follow its precedence constraints and reorders these tasks to follow them, as shown in Eq. 47.
In the first iteration, the local search begins with a rate-monotonic ordered list, which sorts tasks in ascending order of their periods. As discussed in [13], the rate-monotonic ordered list is the desirable starting point since it processes the most constrained tasks (i.e., tasks with the most occurrences) at the start of the procedure when the schedule is mostly empty.
Finally, this local search procedure could easily be used with simulated annealing, tabu search, or another metaheuristic scheme. However, based on the results of [35] and limited experimentation, we utilize the pure local search and try to alleviate its possible disadvantages using the CP approaches described in the following sections.
@@ 4.2 Warm-Start by Constraint Programming
The local search generally works very well. However, with the utilization of instances close to 100%, the approach struggles to find any feasible solution. In that case, it is beneficial to provide the local search with an alternative initial ordered list, which the heuristic maps to a feasible schedule immediately.
@@ 4.2.1 Feasible Schedule for a Single Resource
We use the bin-packing CP formulation for monoprocessor (as described in [13]) based on the height-divisible packing of Section 2.4.3. The CP solver processes one resource at a time and creates a feasible schedule for each. Afterward, these feasible schedules are used to infer a suitable ordered list for the local search.
A set of tasks $\mathcal{J}_{m}$ (with periods $\tau(1),\dots,\tau(j),\dots,\tau(\rho)$) induces an associated set of rectangles $\mathcal{R}_{m}$, whose heights and widths are defined in Section 2.4.3. The set of available heights corresponds to the set of periods of the tasks: $h_{j}=\tau(\rho)/\tau(j)$. Thus, $h_{1}=H$, which is equal to the height of the entire bin, and $h_{\rho}=1$. In Fig. 6, we see that there are two rectangles that span the entire height of the bin and three rectangles of height 1. These are induced by single-resource PSP outlined in Fig. 3(a).
@@
The formulation packs the rectangles of $\mathcal{R}_{m}$ in a height-divisible manner to produce a packing in canonical form. This means that for each rectangle, it holds that all the rectangles packed to its left have a height greater or equal. Fig. 6 shows such packing (unlike in Fig. 3(b)). It was proven in [13]) that if there is a feasible packing, then there exists a feasible canonical packing.
The T-like bold lines in the figure show how the bin is recursively partitioned into sub-bins, where rectangles with smaller heights can be packed, given the height-divisible property of packing. Rectangles with height $h_{1}=H=6$ can only be packed into a single sub-bin that spans the entire figure. The rectangles with height $h_{2}=3$ have two available sub-bins. Finally, the rectangles with $h_{3}=1$ can be packed into six sub-bins. However, since only three of these sub-bin of height 1 are filled with rectangles, the other three have width 0 and are not shown in Fig. 6.
To utilize this sub-bin view of the packing, we partition the rectangles to lists $\mathcal{R}_{m}^{j}$ according to their height $h_{j}$. Each such list is associated with a vector of widths of the rectangles $\mathbf{l}^{j}$ and a vector of integer assignment variables $\mathbf{a}^{{j}}$ with domain $\{0,\cdots,(H/h_{j})-1\}$. The assignment variable determines which sub-bin the rectangle will be packed into. In Fig. 6, height $h_{2}=3$ is relevant for rectangles $\mathcal{R}_{m}^{2}=(3,4,5,6,7)$, with widths $\mathbf{l}^{2}=(17,9,6,5,2)$. The top sub-bin for $h_{2}=3$ is occupied by rectangles 3, 4, and 7. The bottom sub-bin for $h_{2}=3$ is occupied by rectangles 5 and 6. Therefore, the vector of assignment variables is $\mathbf{a}^{{2}}=(1,1,0,0,1)$.
Furthermore, we create a vector of integer variables $\mathbf{L}^{j}=(L^{j}_{0},\dots,L^{j}_{k},\dots,L^{j}_{(H/h_{j})-1})$ for each height $h_{j}$. This vector contains a single variable for each available sub-bin $k$ of height $h_{j}$. This load variable keeps track of how much the sub-bin is filled. For example, the load variable $L^{1}_{0}$, associated with the only sub-bin of height $h_{1}=6$, is set to 12 in Fig. 6. The sub-bins for $h_{2}=3$ are filled $L^{2}_{0}=11$, $L^{2}_{1}=28$ respectively. The width of the bin $w$, into which we pack the rectangles, is given by the shortest period $\tau(1)$.
@@
The well-known global pack constraint [45] (Eq. 49) forces the model to assign to each individual variable of $\mathbf{a}^{{j}}$ an index $k$ of one of the sub-bins with load $L^{j}_{k}$. It also sets the value of $L^{j}_{k}$ to the sum of the widths of assigned rectangles, thus solving a bin-packing subproblem for each distinct height $h_{j}$.
The second constraint Eq. 50 enforces that for each row of height 1, the sum of loads of sub-bins that cross the row (across the different heights) does not exceed the bin’s width. In case of row 2 in Fig. 6, we have:
$L^{1}_{0}+L^{2}_{0}+L^{3}_{2}=12+11+13=36\leq 40=w$ (51)
Once the CP solver finds a feasible HD2D packing in canonical form (given the values of assignment variables $\mathbf{a}^{j}$), the packing is transformed to the schedule of original PSP according to the problems’ equivalence, as was outlined in Section 2.4.3. For details, see [13, 12]. The coordinates of the rectangles determine the start times of the associated tasks.
@@ 4.2.2 Inferring Ordered List from Schedule
An ordered list that should produce a feasible solution for the leftmost heuristic is obtained simply by sorting the tasks by their start time, concatenating the results obtained for all the resources. However, to be used with the predecessor method, this relatively arbitrary ordered list may lead to unexpected behavior since it does not respect inter-resource coupling via precedences. If two consecutive tasks of the chain $J_{i}^{k},J_{i+1}^{k}$ are put in the ordered list one after another, and their independently scheduled start times cause them to collide, the predecessor heuristic will schedule the latter task slightly later to account for it. However, this disrupts the schedule obtained by CP (creating holes) and may eventually lead to infeasibility.
We try to resolve this by employing an order-determining CP model for the predecessor method. In this model, each task $J_{i}^{k}$ of the instance is associated with an integer variable $a^{k}_{i}$. Given the start times of the individual tasks established in the previous section, the CP model is defined using:
$\mathtt{AllDiff}(\left\{a^{k}_{i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \forall J_{i}^{k}\right\})$ (52)
$a^{k}_{i}<a^{l}_{j}$ $\forall m\in\left\{1,\dots,M\right\},\leavevmode\nobreak\ \forall J_{i}^{k}\in\mathcal{J}_{m},$
$\forall J_{j}^{l}\in\mathcal{J}_{m}:t_{i}^{k}<t_{j}^{l}$ (53)
$a^{k}_{i+1}<a^{k}_{i}$ $\forall k\in\left\{1,\dots,\mu\right\},\leavevmode\nobreak\ \forall i\in\left\{1,\dots,n^{k}-1\right\}:$
$t_{i}^{k}+p_{i}^{k}>t_{i+1}^{k}$ (54)
All the tasks across all the resources are processed simultaneously. The solver directly determines the ordered list of tasks with the all-different global constraint Eq. 53 - the value of $a_{i}^{k}$ is the position of $J_{i}^{k}$ in the ordered list. Furthermore, when the successor task $J_{i+1}^{k}$ would be scheduled later than planned due to the start time of its predecessor (creating a hole), constraint Eq. 54 ensures it is processed earlier than its predecessor from the ordered list and the heuristic does not unexpectedly postpone it.
@@ 4.2.3 Feasibility Issues
The aforementioned CP approach consistently finds feasible solutions for most of the 100%-utilized instances; we observed this during the experimentation reported in the following sections of the paper. Nevertheless, when utilization drops, it may happen that the heuristic cannot transform the provided order into a feasible solution. As results of Section 5.1 show, this happens at worst for 10% of instances on the most affected instance set. Furthermore, the local search successfully finds feasible solutions for less-utilized instances even without a warm start.
@@
Figure 7: Schedules documenting the issue with a warm start on less-utilized resources. Notice the gap caused by scheduling task 123 at an incorrect time caused by FFS, leading to infeasibility - task 187 cannot be scheduled.
This phenomenon is shown in Fig. 7. In Fig. 7(a), the schedule of one specific resource produced by the CP warm start is clearly feasible. However, when the FFS heuristic processes tasks row by row, we can see in Fig. 7(b) that it misplaces task 123; this CP model would put the task into the second row, but the heuristic schedules it in the first one, as it operates in a first fit manner, and there is excessive space available. This creates a gap due to the periodicity of the task. This inefficiency in the schedule eventually leads to the state shown in Fig. 7(b). Although we were able to schedule all the tasks until that point, once a task 187 (row 12 in Fig. 7(a)) is to be processed, there is no longer enough space to schedule it successfully.
@@ 4.3 Constraint Programming Matheuristic
The local search performs generally well, but it sometimes struggles to minimize the degeneracy of highly-utilized instances. We demonstrate this using Fig. 8 (specific results are discussed in Sections 5.1 and 5.2). In Fig. 8(a), each curve shows the degeneracy of the best-found solution of each instance as the local search progresses. We see that the local search did not converge for several instances in a given time. However, in other cases, the local search rapidly decreased the degeneracy and stagnated with non-zero degeneracy for the rest of the experiment.
We investigated how the degeneracy is distributed within these stagnating instances. We focused on chains (and tasks) with the longest period $\tau(\rho)$ only. Blue dots in Fig. 8(b) show the values of $\phi^{\tau(\rho)}$. It is the ratio of degeneracy caused by the longest period chains and the total degeneracy for the final solution of each instance (instances are sorted in ascending order of $\phi^{\tau(\rho)}$). We observed that these longest-period chains produce 40%-80% of degeneracy, even though they have just a single occurrence in the schedule’s hyperperiod. Furthermore, orange dots show values of $\Omega^{\tau(\rho)}$, which is the percentage of the utilization allocated to the longest-period tasks on the most utilized resource of an instance. We can see that this value is lower than $\phi^{\tau(\rho)}$. This suggests that while the longest-period tasks do not necessarily utilize the instance the most, their chains significantly contribute to the overall degeneracy. The culprit is probably the fact that the ordered list generated by the local search is generally almost rate-monotonic. Since the FFS heuristic processes these tasks later, there are fewer options to schedule them with low degeneracy.
(a) Degeneracy of the best-found solution over time. Each line corresponds to a single instance.
(b) Ratio of degeneracy $\phi^{\tau(\rho)}$ and ratio of utilization $\Omega^{\tau(\rho)}$ caused by the longest-period chains and tasks.
@@
This observation motivated the use of a matheuristic to improve the schedule further. Similarly, as in [46, 47], the focus is a natural way of decomposition: individual periods. If all the tasks with periods shorter than $\tau(\rho)$ would be fixed, the longest period tasks could be re-scheduled within the complete hyperperiod $\left[0;\tau(\rho)-1\right]$. Other tasks behave as fixed obstacles. Similarly, we can focus on optimizing chains with arbitrary period $\tau(\beta)$.
We focus on chains with $T^{k}=\tau(\beta)$. Let $\mathcal{J}_{m}^{O}$ be a set of tasks with period $\tau(\beta)$ on resource $m$. Let $\mathcal{J}_{m}^{F}$ be a set of all occurrences of other tasks on resource $m$. To avoid any possible collision, tasks with shorter periods are “multiplied” in $\mathcal{J}_{m}^{F}$ to account for each of their occurrences within $\tau(\beta)$; task with a period $\tau(\beta^{\prime})$ adds $\tau(\beta)/\tau(\beta^{\prime})$ properly distanced occurrences to $\mathcal{J}_{m}^{F}$. Tasks with longer periods are simply projected using the modulo operator. If two occurrences in $\mathcal{J}_{m}^{F}$ immediately follow each other, we can merge them into one to simplify the model. These sets are created for all resources.
@@
As in Section 2.4, each resource’s start time is shifted so that a task with the shortest period starts at time 0, let $o_{m}$ be such offset. Let $s_{i}^{k}$ be the shifted core schedule of the task, as it is in Section 2.4. Finally, the set of chains with period $T^{k}=\tau(\beta)$ is denoted as ${C}^{\tau(\beta)}$.
@@
We formulate the following matheuristic CP model. The model uses standard interval variables which tie together start times and processing times and which enable the use of efficient scheduling constraints to handle both the optimized tasks and the fixed occurrences of other tasks. The model tries to re-schedule tasks with period $\tau(\beta)$ to decrease the objective function while keeping the other tasks (their occurrences) fixed:
$\min\leavevmode\nobreak\ \sum_{\mathcal{C}^{k}\in{C}^{\tau(\beta)}}(\lceil S^{k}/\tau(\beta)\rceil-1)$ (55)
$\mathtt{NoOverlap}(\mathcal{J}_{m}^{O}\cup\mathcal{J}_{m}^{F})$ $\forall{m}\in\left\{1,\dots,M\right\}$ (56)
@@
$\mathtt{StartOf}(J_{i}^{k})=s_{i}^{k}$ $\forall{m}\in\left\{1,\dots,M\right\},\leavevmode\nobreak\ \forall J_{i}^{k}\in\mathcal{J}_{m}^{F}$ (57)
@@
$\mathtt{StartOf}(J_{i}^{k})\geq 0$ $\forall{m}\in\left\{1,\dots,M\right\},\leavevmode\nobreak\ \forall J_{i}^{k}\in\mathcal{J}_{m}^{O}$ (58)